ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 177, 2010

Aizenbud, A. and D. Gourevitch The De-Rham theorem and Shapiro lemma for Schwartz functions on Nash manifolds

155

Badzioch, B. and W. Dorabiała A note on localizations of mapping spaces

441

Bass, R. F. and K. Gröchenig Random sampling of bandlimited functions

1

Burness, T. C., E. A. O’Brien and R. A. Wilson Base sizes for sporadic simple groups

307

Colombo, F., I. Sabadini and D. C. Struppa An extension theorem for slice monogenic functions and some of its consequences

369

Deville, R. and J. Rodríguez Integration in Hilbert generated Banach spaces

285

Dorabiała, W. See B. Badzioch Glöckner, H., R. Gramlich and T. Hartnick Final group topolo-

gies, Kac–Moody groups and Pontryagin duality

49

Gorkin, P. and R. Mortini Two new characterizations of Carleson– Newman Blaschke products

267

Gourevitch, D. See A. Aizenbud Gröchenig, K. See R. F. Bass Gramlich, R. See H. Glöckner Grynkiewicz, D. J. On extending Pollard’s Theorem for

t-representable sums

413

Hájek, P. and V. Montesinos Boundedness of biorthogonal systems in Banach spaces

145

Haïly, A., A. Kaidi and A. Rodríguez Palacios Algebra descent spectrum of operators

349

Hartnick, T. See H. Glöckner Henniart, G. and B. Lemaire Pseudo-coefficients des séries

κ-discrètes de GL(n, F)

189

Hernández Cifre, M. A. and E. Saorín On inner parallel bodies and quermassintegrals

29

Hurtado, A. Instability of Hopf vector fields on Lorentzian Berger spheres

103

Johansson, A., T. M. Jordan, A. Öberg and M. Pollicott Multifractal analysis of non-uniformly hyperbolic systems 125 Jordan, T. M. See A. Johansson Kaidi, A. See A. Haïly Kaminski, J. Y., A. Kanel-Belov and M. Teicher Multi-Secant

lemma

253

Kanel-Belov, A. See J. Y. Kaminski Kang, M.-H., W.-C. W. Li and C. Wang The zeta functions of

complexes from PGL(3): a representation-theoretic approach

335

Lemaire, B. See G. Henniart Li, W.-C. W. See M.-H. Kang Montesinos, V. See P. Hájek Mortini, R. See P. Gorkin O’Brien, E. A. See T. C. Burness Öberg, A. See A. Johansson Pollicott, M. See A. Johansson Raja, C. R. E. and R. Shah Distal actions and shifted convolution

property

391

Rodríguez, J. See R. Deville Rodríguez Palacios, A. See A. Haïly Sabadini, I. See F. Colombo Saorín, E. See M. A. Hernández Cifre Schlage-Puchta, J.-C. The subgroup growth spectrum of virtually

free groups

229

Shah, R. See C. R. E. Raja Struppa, D. C. See F. Colombo Teicher, M. See J. Y. Kaminski Wang, C. See M.-H. Kang Wilson, R. A. See T. C. Burness